OpenStudy (anonymous):

obtain the general solution: (1-x)y'=y^2 the ans. is 1= ylnC(1-X) but when i answered it, it was 1 = -ylnC(1-x) how can i remove then negative(-) sign?

OpenStudy (tkhunny):

You got the sign right on the y-side, but you forgot your chain rule on the x-side. BTW \(\int \dfrac{1}{1-x}\;dx = -ln|1-x| + C \ne -ln(1-x) + C\)

OpenStudy (anonymous):

uh, whats the difference b/w them?

OpenStudy (tkhunny):

Absolute values and not absolute values? For the logarithm function, you must KNOW the argument is positive.

OpenStudy (anonymous):

oh, so to make the natural logarithm positive, you make (1-x) as absolute values therefore |1-x| right?

OpenStudy (tkhunny):

This is a common point of confusion. \(\dfrac{d}{dx}ln(x) = \dfrac{1}{x}\) However, \(\int \dfrac{1}{x}\;dx = ln|x| + C\) When we did the derivative, we already had a valid logarithm function, so we already know x > 0. When we did the integral, we don't know anything about x so we have to force it to be positive in order to introduce the logarithm. Making a little bit of sense? If you KNOW x is positive, then you can just forget about it.

OpenStudy (anonymous):

oh, thank you!